OpenStudy (anonymous):

Write an equation of the live that passes through the given points. (-3,6) and (1,2) The equation is

OpenStudy (anonymous):

use two point formula

OpenStudy (alekos):

just use the point slope form y-y1=m(x-x1) where m= (y2-y1)/(x2-x1)

OpenStudy (anonymous):